OpenStudy (anonymous):

what are the coordinates of the vertex of the graph of the function y = -x^2 + 6x - 11? a. (3,-2) b. (3,16) c. (-3,-29) d. (-3,-20)

OpenStudy (freckles):

so \[-x^2+6x-11 \\ =-(x^2-6x)-11 \\ =-(x^2-6x+?)-11+?\] now what does that one question need to be in order to complete the square

OpenStudy (freckles):

hint: \[x^2+kx+?=(x+\frac{k}{2})^2 \\ \text{ the } ? \text{ needs to be } (\frac{k}{2})^2\]

OpenStudy (freckles):

so what is you k here?